OpenStudy (anonymous):

while finding gcf using euclid algorithm, why do we chase after remainder and quotient after each step ? dont get why gcf of two numbers equal gcf of quotient and remainder. could somebody care to explain plzz ?

OpenStudy (anonymous):

We try to find gcf of 2445 and 652. We have ``` step1 : 2445 = 652*3 + 489 step2 : 652 = 489*1 + 163 step3 : 489 = 163*3 + 0 ``` thus the gcf of 2445 and 652 is 163 :S i dont get why gcf(2445, 652) equals gcf(652, 489)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

okay thank you :)

jimthompson5910 (jim_thompson5910):

I think I found a page that explains it, but I'm going to try and translate it into something that makes more sense, one more sec (sorry for the long wait)

OpenStudy (anonymous):

May I answer, @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

sure go for it

OpenStudy (anonymous):

Gratzie. All right, so, let's say we have a number \(d\) that divides your two numbers \(a, b\). If \(d\) divides \(a\), then we can multiply \(d\) by some number (let's call it \(d_a\)) to equal \(a\), similarly we do the same for \(b\) and call this number \(d_b\). So we know: \[ d\cdot d_a=a\\ d\cdot d_b=b \]So, if we have a linear combination of a and b, expressed as: \[ ax+by \]For some integers \(x, y\), then, we can say: \[ ax+by=d\cdot d_a\cdot x+d\cdot d_b\cdot y=d(d_a x+d_b y) \]Which means that d also divides any linear combination of these two numbers. Thus, since the remainder is a linear combination of the two numbers, the GCD also divides the remainder, as well as the two numbers. Another way to show this is to give Bezout's Identity.

OpenStudy (anonymous):

@LolWolf thanks! so remainder(489) is divisible by gcf i get that. cuz 489 = 2445-652*3 which is a combination of the two input numbers. how we conclude gcf of remainder and the divisor gives the gcf of input numbers from this ?

OpenStudy (anonymous):

All right, so we know that the GCD (let's call it \(g\)) must divide the pen-ultimate remainder, \(r\), right? (due to the above proof) So, therefore, we know that all divisors of both numbers divide \(r\), and, thus, g being one of those numbers, also divides \(r\), which implies that \(g\leq r\). We also know that \(r\) divides a linear combination of the two numbers (since \(r\) also divides each number), so, \(r\) must also divide \(g\), which implies that \(r\leq g\). Since: \[ r\leq g\wedge g \leq r\implies g=r \]Thus we are done. QED.

OpenStudy (anonymous):

LolWolf im not able to follow fully.. So, therefore, we know that "all divisors of both numbers divide r," and, thus, g being one of those numbers, also divides r, which implies that g≤r. there you mean "all divisors common to both numbers divide r" is it ?

ganeshie8 (ganeshie8):

id like to understand this too :p

OpenStudy (anonymous):

How you made that grey bracket? Can you tell me ??

OpenStudy (anonymous):

Grey box sorry..

OpenStudy (anonymous):

tick tick tick tick tick tick

OpenStudy (anonymous):

``` ```

OpenStudy (anonymous):

\`\`\` \`\`\`

OpenStudy (anonymous):

This you have made I know the code for that..

OpenStudy (anonymous):

Sorry I did not get,,

OpenStudy (anonymous):

\`\`\` \`\`\`

OpenStudy (anonymous):

write anything in between 3 back ticks.

OpenStudy (anonymous):

for getting greybox...

OpenStudy (anonymous):

How we make back ticks ??

ganeshie8 (ganeshie8):

you have the key on your keybord "~"

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

Oh yeah I saw that..

OpenStudy (anonymous):

` ` ` ` ` `

ganeshie8 (ganeshie8):

water good to see u u maintaining low profile these days :)

OpenStudy (anonymous):

Low profile meaning ???

ganeshie8 (ganeshie8):

means not seeing u ol much

OpenStudy (anonymous):

Oh my college has been started so I don't get time to be online here..

OpenStudy (anonymous):

could you guys help me plzz

OpenStudy (lgbasallote):

psh who needs college

OpenStudy (anonymous):

What exactly is your question Sara ??

OpenStudy (anonymous):

my first reply just after the question i dont get why gcf(2445, 652) equals gcf(652, 489)

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Can't you try by working them out??

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Can you find gcf of 2445 and 652 ??

OpenStudy (anonymous):

it works. does not look obvious

OpenStudy (anonymous):

i can find using both euclid algorithm and prime factors

OpenStudy (anonymous):

Then use Euclid Algorithm..

OpenStudy (anonymous):

i get 163. how i got it i have no idea while using euclid algorithm where as prime factors is quite obvious

OpenStudy (anonymous):

Now find the same for 652 and 489..

OpenStudy (anonymous):

using prime factors ?

OpenStudy (anonymous):

You can use both,, But use Euclid Algorithm here too..

OpenStudy (anonymous):

step1 : 2445 = 652*3 + 489

OpenStudy (anonymous):

step2 : 652 = 489*1 + 163

OpenStudy (anonymous):

Go ahead...

OpenStudy (anonymous):

i am simply copying my first reply and pasting here...

OpenStudy (anonymous):

Yes I know but go for second one now..

OpenStudy (anonymous):

step2 ?

OpenStudy (anonymous):

`652 = 489*1 + 163 ` ` ` 489 = 163*1 + etc etc` ` Like this..

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

163*2 I think..

OpenStudy (anonymous):

*3

OpenStudy (anonymous):

step3 : 489 = 163*3 + 0

OpenStudy (anonymous):

Yep.. So again you got 163 don't you think the gcf of both are equal that is 163 ??

OpenStudy (anonymous):

i just dont see how the logic is flowing.. . step1 to step2, why we took divisor and remainder ?

OpenStudy (anonymous):

step3 : 489 = 163*3 + 0 i think, gcf of 489 and 163 is 163

OpenStudy (anonymous):

im sorry just trying to understand

OpenStudy (anonymous):

According to Euclid Algorithm, We just divide two numbers: After that the remainder got in first step becomes quotient and the quotient of first step becomes dividend again divide them.. and same process goes on until you get 0..

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i dont understand the second step

OpenStudy (anonymous):

See, gcf of 652 and 489 is 163.. and gcf of 489 and 163 is also same..

OpenStudy (anonymous):

Second step where??

OpenStudy (anonymous):

Basically what do you mean by gcf Sara can you tell me ??

OpenStudy (anonymous):

greatest common factor

OpenStudy (anonymous):

I think it is just the full form of gcf.. I am asking about its meaning..

OpenStudy (anonymous):

greatest integer that goes in both a and b

OpenStudy (anonymous):

the gcf of a and b

OpenStudy (anonymous):

I tell you.. Basically, it means the greatest factor that will be common to both and will divide both simultaneously.. For example: For 3 and 9, what factor can divide both ??

OpenStudy (anonymous):

3 is the factor that is common to both 3 and 9 and will divide 3 and 9 both..

OpenStudy (anonymous):

3 = 3^1 9 = 3^2 gcf = 3^1

OpenStudy (anonymous):

finding gcf using prime factors is obvious. i get it

OpenStudy (anonymous):

and 3 here is greatest factor too.. Because you cannot have factor greater than 3 like 4 and 5 that can divide 3 and 9 both at same time.. Getting Sara??

OpenStudy (anonymous):

its euclif algorithm that makes me insane

OpenStudy (anonymous):

wit plz

OpenStudy (anonymous):

See, prime factorization and Euclid are just the way to find gcf but you must understand first what is gcf..

OpenStudy (anonymous):

wait*

OpenStudy (anonymous):

Take your time...

OpenStudy (anonymous):

Because you cannot have factor greater than 3 like 4 and 5 that can divide 3 and 9 both at same time.. i get this

OpenStudy (anonymous):

yes that is the greatest and common and factor too that is why that is called gcf..

OpenStudy (anonymous):

thats the reason in prime factors we search for greatest number of common factors..

OpenStudy (anonymous):

yeaa

OpenStudy (anonymous):

Oh so basically you are asking the working of Euclid Algorithm..

OpenStudy (anonymous):

yea just this : :S i dont get why gcf(2445, 652) equals gcf(652, 489)

OpenStudy (anonymous):

See if 163 is the greatest common factor then that will divide 2445, 652 and 489 at the same time.. On dividing by 163 at the same time you will get : 15, 3 and 1..

OpenStudy (anonymous):

It is same as; As table of 3 contains 9 15 27 45 etc Similarly table of 163 contains: 2445, 652 and 489..

OpenStudy (anonymous):

Sorry, I'm a little late to the party (I have to sleep!). But, the point of my first proof was to show that the pen-ultimate remainder (i.e. the one before the last remainder, which equals zero) \(r_p\) is also, in fact a linear combination of the two numbers, right? Because if we rearrange division algorithm, we get, for some number and quotient \(n, q\) and divisor and remainder \(d, r\): \[ n=qd+r\implies\\ r=n-qd \]Here, \(r\) is also a linear combination of both \(d, n\). This means, by our previous proof that \(\text{GCD}(n,d)=\text{GCD}(r,d)\). And that the GCD divides \(r\) (and therefore \(r_p\)), thus \(g\le r_p\). But, since the pen-ultimate remainder must have divided it's preceding number (due to the last step equaling zero), it therefore must divide the GCD (which we define as: "the divisor of two numbers \(a,b\) such that all other common divisors divide it), if it divided a linear combination of it. By this, then, we have that \(r_p\le g\). Putting these two conditions together, we get that \(r_p=g\).